Dan is gliding on his skateboard at. 4, 0m/s. He suddenly jumps backwards off his skateboard, kicking the skateboard forward at 8,0 m/s. How fast is Dan going as his feet hit the ground? Dan's mass is 50kg and the skateboards mass is 5,0kg.
1
Expert's answer
2020-05-04T12:51:04-0400

As per the given question,

initial speed of the skateboard"(u)=4m\/sec"

final speed of the Dan "(v)=8.0 m\/sec"

Dan's mass (M)=50kg

skateboards mass (m) = 5.0kg

Let the final speed of the skateboards =V

Applying the conservation of momentum,

"\Rightarrow (m+M)u=-M\times 8+m\times V"

"\Rightarrow V=\dfrac{(m+M)u}{m}+\dfrac{M\times 8}{m}"

"\Rightarrow V=\dfrac{55\times 4+8\times 5}{5}=52 m/sec"